Computer Science 61A Course Reader Spring, 2003

Table of Contents

Semester Schedule  2
Ergonomic Information  3
Weekly Assignments  7
SICP Errata  30

Projects:

Project 1 (Twenty-one)  31

Project 2 is in the textbook! (Section 2.2.4)

Project 3 (Adventure game)  35
Project 4 (Logo interpreter)  43

Old Exams: Note: Prior to Spring, 2003, Midterm 1 covered through 2.2.1; Midterm 2 covered through OOP; Midterm 3 all of Chapter 3.

Fall 2001 exams and solutions  55
Fall 2000 exams and solutions  95
Fall 1999 exams and solutions  138

Lab assignments  colored pages

Reference Documents:

Object-Oriented Programming — Above the line view  1
Reference Manual for the OOP Language  11
Object-Oriented Programming — Below the line view  14
Highlights of GNU Emacs  21
Emacs Quick Reference Guide  41
An Investigation of the Therac-25 Accidents  45
Revised^{5} Report on Scheme  71
Lecture Notes  121